2017/3/6(Tue.)

 * Remote participants by eZuce: Kimura

 * Report of work (*:responsible person)

  * Alignment: Enomoto
   * Beam was found again at PD and Tcam of ETMX and Tcam of ETMY.

  * BS: Mark
   * We lifted up the optical table.

  * IOO:
   * IMMT2 cannot be seen by camera. We will work on camera this week.

  * CRY: mirror temperature ~40K.

 * Work plan (*:responsible person)

  * Alignment: Enomoto
   * We will find beam at the ETMY or ETMX by dithering PR3.

  * AOS: Akutsu, Michimura
   * Items PR2 chamber will be replaced, and green light will be used in the afternoon.
   * WAB work at the IYA.

 * Laser hazard: AM: PSL, PM: Upstream SR2-3 line, and Y-end for both IR and green.

 * Weekly

  * week of 3/5
   * IYC door will be closed.
   * Exchanging items and cleaning inside the PR2 chamber

  * week of 3/12
   * TMS installation at ETMY.
   * Connecting ducts at the corner area.

  * week of 3/19
   * Vacuum test at at the corner area.
   * TMS alignment using green light at ETMY.
   * Beam alignment to REFL port.


 * Issues